Find the uncertainty in a calculated electrical potential difference from the measurements of current and resistance. Electric potential difference depends on current and resistance according to this function V(I,R) = IR. Your measured current and resistance have the following values and uncertainties I = 3.3 Amps, delta I space equals space 0.9 Amps and R = 87.5 Ohms and delta R space equals space 0.8 Ohms. What is the uncertainty in the , delta V
Answer #1

Given,

I = 3.3 A; ΔI = 0.9 A; R= 87.5 ohms; ΔR = 0.8 ohms

using ohm's Law, V=IR = 3.3*87.5 = 288.75 V

Now to calculate the error in V

ΔV/V = sqrt((ΔI/I)2 + (ΔR/R)2)

ΔV/288.75 = sqrt((0.9/3.3)2 + (0.8/87.5)2)

ΔV = 78.794 V